a transformer consists of 2000 turn primsry coil and 300 turn secondary with the secondary coil amp is 9. what is the primary current? i get 1.35
Answer #1

Solution:

Using

N_{p}I_{p}=N_{s}I_{s}

\left ( 2000 \right )I_{p}=\left ( 300 \right )\left ( 9A \right )

I_{p}=1.35A